OpenStudy (anonymous):

J K 2 -12 4 -24 6 -36 8 -48

OpenStudy (anonymous):

Which function rule would help me find the values in the table? @tHe_FiZiCx99

OpenStudy (anonymous):

Choices are: k = -12j k = 6j k+ j- 12 k = j-6

OpenStudy (the_fizicx99):

Hm, pick two points for me, I'll guide you!

OpenStudy (anonymous):

That is the problem no idea how to start. confused me when it says -

OpenStudy (anonymous):

But, I will try.

OpenStudy (anonymous):

I think k = -12j

OpenStudy (anonymous):

multiplies

OpenStudy (the_fizicx99):

Go for it! Pick two of these points: {2 -12} {4 -24} {6 -36} {8 -48}

OpenStudy (anonymous):

Did I get it right above?

OpenStudy (the_fizicx99):

We can try and see, pick two points, it's really easy :>

OpenStudy (anonymous):

you mean pick like 2 and -12 ?

OpenStudy (anonymous):

wouldn't that be -24 ?

OpenStudy (the_fizicx99):

Yes!

OpenStudy (the_fizicx99):

No not like that, points such like {4 -24} and {8 -48}

OpenStudy (anonymous):

:) did I get it right with my answer? k -12j

OpenStudy (the_fizicx99):

I haven't done this problem thoroughly, we can do it together and I can tell you if you're doing it correctly. :P Or, you can show me how you got k - 12j :o

OpenStudy (anonymous):

2 x -12 = -24

OpenStudy (the_fizicx99):

Ok, now how did you find J and K ?

OpenStudy (anonymous):

multiplied? or I just divided 2 x -24 = -48

OpenStudy (the_fizicx99):

Hm, just pick two points, you can't find a function rule like that, it won't work :P

OpenStudy (anonymous):

ok. { 2 -12} {4 -24}

OpenStudy (the_fizicx99):

Great, are you familiar with the slope formula?

OpenStudy (anonymous):

ohhhhh it is k = - 6j !!!

OpenStudy (anonymous):

because 6 multiplied by all the j's numbers equals the k right

OpenStudy (the_fizicx99):

Yes, it is k = -6j, but do you know how to arrive at that answer without using the options? :O Gratz for finding it though -48 = -6(8) -48 = -48

OpenStudy (anonymous):

oh ok I am soooo writing that note down. yeah that helps. Thank you.!! Can I ask you later another question if your on line?

OpenStudy (the_fizicx99):

Yup :P

OpenStudy (anonymous):

awesome! Thanks!

OpenStudy (the_fizicx99):

yw